# A Linear Quantum Coupler for Clean Bosonic Control

### v2

@@ VI Conclusions and Outlook
The parity-protection in the LINC could also provide important advantages in its power handling and multi-tone operation. This is useful in multiple contexts, with the simplest being the activation of a simultaneous parametric coupling between multiple neighboring elements. If the LINC were used as an amplifier, an array of LINCs should perform equivalently to an array of balanced RF-SQUIDs, potentially outperforming both SNAIL and regular RF-SQUID based amplifiers. Such an implementation would potentially allow simultaneous gain at multiple frequencies, easing the constraints on multiplexed readout [55, 56]. Finally, the LINC could also simultaneously activate multiple types of parametric processes, giving rise to new bosonic control techniques. For example by activating a resonant beamsplitting and two-mode squeezing between two oscillators in the high-Q regime, one could realize a direct parametric quadrature-quadrature coupling between them, enabling two-qubit gates for the Gottesman-Kitaev-Preskill (GKP) code [57, 58].
The advantages predicted in our simulations will require experimental proof, which is the subject of immediate future work. Specifically, one must solve the engineering challenges of delivering the differential bias and drive in a compact architecture, and fabricating high-Q shunting inductors or junction-array. However, the analysis of the LINC’s sensitivity to small asymmetries shows that its experimental implementation need not be impractically precise, and that the coupler should perform well even with realistic imperfections. Overall, the LINC is a promising new element in the toolbox of superconducting quantum circuits, offering a unique combination of linearity, high-fidelity control, and practical design constraints.
@@ VII Acknowledgements
We are grateful for discussions with P. Winkel on high-kinetic inductance shunting inductors and with S. J. de Graaf on the SNAIL as a three-wave mixer. We also thank M. Devoret and C. Zhou for helpful discussions about the LINC Hamiltonian and dephasing. We thank M. Hayes and W. Oliver for general discussions about the coupler.
This research was sponsored by the Army Research Office (ARO) under grant no. W911NF-23-1-0051, and by the U.S. Department of Energy (DoE), Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage ($\text{C}^{\text{2}}$) under contract number DE-SC0012704.The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the ARO, DoE or the US Government. The US Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. L.F. and R.J.S. are founders and shareholders of Quantum Circuits Inc. (QCI).
@@ Appendix A Analyzing the LINC circuit
For the rest of this section, let us assume that all symmetry constraints in the circuit are satisfied, specifically that the outer junctions have equal Josephson energy $E_{J}$, and any DC flux and drive enter both loops symmetrically. We will explore the effects of deviating from this ideal in appendix C. This constraint lets us set $\hat{\phi}_{\text{asym}}=0$, thus giving:
@@
$\begin{gathered}g_{mn}[\phi_{\text{DC}}]:=\dfrac{1}{m!\>n!}\sum_{k=0}^{n}\binom{n}{k}\left.u_{m,n-k}\;\frac{\partial^{k}\theta_{zpf}^{m}}{\partial\phi_{d}^{k}}\right|_{\phi_{d}=\phi_{\text{DC}}},\\ u_{m,k}:=\left.\dfrac{\partial^{m+k}\;U_{\text{LINC}}}{\partial\hat{\theta}^{m}_{c}\;\partial\phi_{d}^{k}}\right|_{\hat{\theta}_{c}=0}.\end{gathered}$ (22)
@@
Note that this is almost exactly half the flux noise sensitivity $d\omega_{L}/d\phi_{\text{DC}}=4g_{2,1}+d\alpha_{L}/d\phi_{\text{DC}}$, due to conventions for the definition of $g_{3wm}$ chosen in eq. 10. This analytic formula matches the numerically derived three-wave mixing strength reasonably well (LABEL:appx_fig1b). The latter is calculated from a Floquet simulation of either a self-squeezing of the LINC (with an added Kerr to induce $|0\rangle\leftrightarrow|2\rangle$ oscillations), or a full beamsplitting process (see appendix B). It is also simple to derive the relative sensitivity of this process to flux noise, for eq. 6 in the main text:
@@ Appendix B Driven operations with the LINC
Figure 6: Driven coupler infidelity a, Dispersion in beamsplitting strength as a function of coupler (undriven) state $|0\rangle,\dots,|3\rangle$. The LINC (grey) has significantly less dispersion than the SNAIL (yellow), which is especially important since the SNAIL is always driven to a displaced state and has coherent shot noise, but the LINC remains in its undriven ground state. b, Dispersion in the beamsplitting resonance condition as a function of coupler state. The LINC practically has no dispersion at low drive amplitudes, while the SNAIL has an idle dispersive shift to the resonators even when its self-Kerr is set to zero. At larger drives strengths these dispersions become similar. c, Steady-state impurity of the driven LINC and the (charge) driven SNAIL, in the presence of decay and flux-noise dephasing, as a function of drive frequency and amplitude. This represents the available drive space for either three-wave mixer without being driven to a mixed state. Any steady state impurity can spoil the parametric process fidelity due to the above state-dependent shifts.
fig]appx_fig3
To numerically demonstrate such a parametric beamsplitting process, and highlight the difference between the LINC and the SNAIL coupler devices, we consider the specific frequency stack of $\omega_{a}=2\pi\times 4.9$GHz, $\omega_{b}=2\pi\times 6.0$GHz, and $\omega_{c}=2\pi\times 6.5$GHz, with Rabi coupling strengths of $g_{ac}=2\pi\times 120$MHz and $g_{bc}=2\pi\times 50$MHz. We choose the LINC circuit parameters as $E_{C}=2\pi\times 100$MHz, $E_{J}=2\pi\times 15.84$GHz, and $E_{L}=2\pi\times 52.8$GHz. With M LINC loops in an array, the system Hamiltonian is given by
@@
For comparison on driven performance, we also simulate a SNAIL circuit with parameters $E_{C}=2\pi\times 100$MHz, $E_{J}=2\pi\times 276$GHz, and $\alpha=0.193$. We model the driven SNAIL Hamiltonian in the displaced frame as
@@
where $M$ is the number of SNAIL loops in an array, $N$ is the number of Josephson junctions in the array of each SNAIL loop, and $\theta_{\text{eq,i}}$ is the frustration phase across the i-th branch of the SNAIL loop, as a result of the dc flux penetrating the loop. For consistency, we denoted the phase displacement of the SNAIL as $\phi_{AC}$, even though it has a different physical origination than flux modulation in the LINC circuit. Under these parameters, the SNAIL circuit yields the same frequency of $\omega_{c}=2\pi\times 6.5$GHz as the LINC, as well as the same $g_{3wm}$, at its Kerr-free point of $\Phi_{\text{ext}}=0.442\Phi_{0}$.
@@
We employ time-dependent perturbation theory for both couplers in the Floquet mode basis [26], to calculate the Alice-Bob beamsplitter rate given a particular coupler state, $|m\rangle$. This amounts to finding the coefficient of the effective Hamiltonian,
[…]
@@
When the coupler is in its ground state ($m=0$), this calculation provides the beamsplitter rate reported in the main text. When the coupler is in an excited state, we study the state-dependent shifts of the beamsplitter rate and resonance condition (LABEL:appx_fig3a,b), which can dephase the beamsplitting process if the coupler is spuriously excited by the parametric drive. Specifically, the amount of state-dependent dispersion in $g_{\text{BS}}$ and $\Delta_{\text{BS}}$ sets the magnitude of process infidelity, given a driven incoherent excitation to that coupler state.
Finally, using the Floquet-Markov method, we can calculate the transition rates among different Floquet modes, their time-domain evolution, as well as the steady state of the system under drive [62, 63]. We assume a thermal relaxation spectrum of the coupler to be $\Gamma_{1}(\omega)=\frac{\omega}{\omega_{q}}(1+n_{B}(\omega))\gamma_{1}$, where $\gamma_{1}[\omega]=(26.7\mu\text{s})^{-1}$, and $n_{B}(\omega)=1/(e^{\hbar\omega/k_{B}T}-1)$ is the Bose occupation factor. We also assume $1/f$ dephasing spectrum for the coupler, $S_{\Phi\Phi}[\omega]=\frac{(1\mu\Phi_{0})^{2}}{\omega}$, with a cut-off frequency of 1 Hz. We examine the impurity of the steady state over different drive frequencies and amplitudes, as a proxy for process infidelity induced by spurious coupler excitation, as well as other parasitic coupler-resonator interactions that may result in a mixed state. The contrast in the available drive space for the driven LINC and driven SNAIL (see LABEL:appx_fig3c) illustrates the clear advantage of the LINC as a cleaner parametric coupler.
Appendix C Effect of junction and DC flux asymmetries
@@
Figure 7: LINC Kerr in the presence of asymmetries a, Kerr of the LINC mode at $\phi_{\text{DC}}=\pi/2$, computed by direct diagonalization. Black contour lines correspond to $|\alpha_{L}|=250,500,750,\text{and }1000$ KHz respectively. b, Operating point $\phi_{d}$ where the LINC is Kerr-free, as a function of asymmetry. With a minor change in the symmetric DC flux, the Kerr induced by small asymmetry can be nulled. Black contour lines correspond to flux values from $298\%\times\pi/$ 2 to $2102\%\times\pi/$ 2 in $0.5\%$ increments around the asymmetry-free DC operating point. c, third-order nonlinearity of the LINC dipole mode, $g_{30}$. With finite junction asymmetry, the self-Kerr and cross-Kerr-free points will differ. Contour lines correspond to $|g_{30}|=3,6,9,\text{and }12$ MHz respectively.
@@
$\begin{split}U/E_{L}=&\frac{1}{2}(\hat{\theta}_{c}-2\phi_{\Delta}/3)^{2}\\ +\beta_{\Sigma}\sin(\phi_{AC})&\left[\cos(\phi_{\Delta}/3)\cos(\hat{\theta}_{c})-\sin(\phi_{\Delta}/3)\sin(\hat{\theta}_{c}))\right]\\ +\beta_{\Delta}\cos(\phi_{AC})&\left[\cos(\phi_{\Delta}/3)\sin(\hat{\theta}_{c})+\sin(\phi_{\Delta}/3)\cos(\hat{\theta}_{c}))\right].\end{split}$ (42)
@@
Figure 8: Driven process strength in the presence of asymmetries a, Fractional strength of the desired $g_{21}$ process in the presence of asymmetries. We see that the strength is primarily dependent upon the asymmetric flux $\phi_{\Delta}$, and is second order with respect to both asymmetries. b, strength of $g_{11}$, the linear coupling of the LINC mode to a symmetric flux drive, as compared to the LINC’s linear coupling to a charge drive $g_{11}^{\text{charge}}\approx E_{L}\theta_{zpf}$. c, Parity protection within the 3rd order nonlinearity. Strength of $g_{12}\phi_{\text{AC}}^{2}$, that would permit subharmonic-driving-like processes, as compared to the desired $g_{21}\phi_{\text{AC}}$, is shown as a function of asymmetry for the drive value $|\phi_{\text{AC}}|=0.2\pi$. The above simulations are performed by numerically computing appropriate derivatives of the potential and $\theta_{zpf}$ as detailed in eqs. 43, 44 and 45. The behavior of these driven process strengths with respect to asymmetry has been separately verified by floquet and/or time-domain simulations.
@@
Figure 9: Effects of parasitic inductance a, Circuit diagram for the LINC with parasitic loop inductance ($L_{\text{loop}}$) and series inductance ($L_{\text{P}}$). We analyze their effects by quantifying their fractional branch participations $\beta_{l}=L_{\text{loop}}/L_{\text{J}}$ and $\beta_{p}=L_{\text{P}}/L_{\text{S}}$. b, Kerr at the operating point $\phi_{\text{DC}}=\pi/2$ as a function of inductance fractions $\beta_{l}$ and $\beta_{p}$. The series inductance has a minimal effect, but a large loop inductance can introduce significant Kerr. c, Frequency and Kerr as a function of operating point $\phi_{\text{DC}}$, in the presence of parasitic inductances. A series inductor changes the participation of the loop and therefore the frequency tuning range, but does not change the Kerr-free point. A loop inductance has minimal changes to the frequency, but can change the LINC’s Kerr. For reasonable loop sizes ($\leq 100\;\mu$m, $\beta_{p}\sim 0.01$), the Kerr can be nulled by only a small shift in the operating point.
@@
Appendix D Effects of parasitic loop inductance
We now analyze the effect that a loop parasitic inductance has on the static nonlinearity of the LINC potential energy. In particular, we assume each Josephson junction in the outer SQUID loop has a stray series inductance ($L_{\text{P}}$ in LABEL:appx_fig4a), and apply nonlinear current conservation methods developed in [65]. We give intuition for and setup the equations for the analysis below, and then evaluate the equations numerically for actual comparison. Note that because the loop inductance always shares current with an outer junction, it is difficult to directly give analytic solutions that are non-parametric. The case of the series inductance in the coupler arms ($L_{\text{P}}$ in LABEL:appx_fig4a) does not face this issue, and we treat it fully analytically in the next section.

### v3

@@ VI Conclusions and Outlook
The parity-protection in the LINC could also provide important advantages in its power handling and multi-tone operation. This is useful in multiple contexts, with the simplest being the activation of a simultaneous parametric coupling between multiple neighboring elements. If the LINC were used as an amplifier, an array of LINCs should perform equivalently to an array of balanced RF-SQUIDs, potentially outperforming both SNAIL and regular RF-SQUID-based amplifiers. Such an implementation would potentially allow simultaneous amplification of several signals without degradation from intermodulation products [55, 32, 56], easing the constraints on multiplexed readout [56, 57]. Finally, the LINC could also simultaneously activate multiple types of parametric processes, giving rise to new bosonic control techniques. For example by activating a resonant beamsplitting and two-mode squeezing between two oscillators in the high-Q regime, one could realize a direct parametric quadrature-quadrature coupling between them, enabling two-qubit gates for the Gottesman-Kitaev-Preskill (GKP) code [58, 59].
The advantages predicted in our simulations will require experimental proof, which is the subject of immediate future work. Specifically, one must solve the engineering challenges of delivering the differential bias and drive in a compact architecture, and fabricating high-Q shunting inductors or junction-array. Thankfully, the LINC incorporates standard circuit elements and does not operate in an exotic parameter regime. We do not foresee any major fabrication challenges, as the LINC shares the topology and parameter range of the ATS, which has been implemented in multiple experiments [20, 60, 61, 62]. As the LINC must be operated away from any sweet-spot (since $g_{\text{3wm}}\propto\frac{\partial\omega}{\partial\Phi}$), it is vital to mitigate flux noise, but the LINC does not require unreasonably low flux noise to perform well. For scaling up, the LINC is readily compatible with existing processor architectures comprising any mixture of qubits and resonators (bosonic modes), and can act as a parametric coupler between them. As the control lines for the LINC are nearly identical to the DC-SQUID, any architecture capable of fast-flux control could utilize LINCs. While cross-talk between LINCs would be heavily implementation-specific, we note that the platform of flux-tunable transmons has seen great success in large-scale implementations [63, 64]. Additionally, our analysis of the LINC’s sensitivity to small asymmetries shows that its experimental implementation need not be impractically precise, and that the coupler should perform well even with realistic imperfections. Overall, the LINC is a promising new element in the toolbox of superconducting quantum circuits, offering a unique combination of linearity, high-fidelity control, and practical design constraints.
@@ VII Acknowledgements
We are grateful for discussions with P. Winkel on high-kinetic inductance shunting inductors and with S. J. de Graaf on the SNAIL as a three-wave mixer. We also thank M. Devoret and C. Zhou for helpful discussions about the LINC Hamiltonian and dephasing. We thank M. Hays and W. Oliver for general discussions about the coupler.
This research was sponsored by the Army Research Office (ARO) under grant no. W911NF-23-1-0051, and by the U.S. Department of Energy (DoE), Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage ($\text{C}^{\text{2}}$QA) under contract number DE-SC0012704.The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the ARO, DoE or the US Government. The US Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. L.F. and R.J.S. are founders and shareholders of Quantum Circuits Inc. (QCI).
@@ Appendix A Analyzing the LINC circuit
For the rest of this section, let us assume that all symmetry constraints in the circuit are satisfied, specifically that the outer junctions have equal Josephson energy $E_{J}$, and any DC flux and drive enter both loops symmetrically. We will explore the effects of deviating from this ideal in appendix D. This constraint lets us set $\hat{\phi}_{\text{asym}}=0$, thus giving:
@@
$\begin{gathered}g_{mn}[\phi_{\text{DC}}]\mathrel{\mathop{\ordinarycolon}}=\dfrac{1}{m!\>n!}\sum_{k=0}^{n}\binom{n}{k}\left.u_{m,n-k}\;\frac{\partial^{k}\theta_{zpf}^{m}}{\partial\phi_{d}^{k}}\right|_{\phi_{d}=\phi_{\text{DC}}},\\ u_{m,k}\mathrel{\mathop{\ordinarycolon}}=\left.\dfrac{\partial^{m+k}\;U_{\text{LINC}}}{\partial\hat{\theta}^{m}_{c}\;\partial\phi_{d}^{k}}\right|_{\hat{\theta}_{c}=0}.\end{gathered}$ (22)
@@
Note that this is almost exactly half the flux-noise sensitivity $d\omega_{L}/d\phi_{\text{DC}}=4g_{2,1}+d\alpha_{L}/d\phi_{\text{DC}}$, due to conventions for the definition of $g_{3wm}$ chosen in eq. 10. This analytic formula matches the numerically derived three-wave mixing strength reasonably well (LABEL:appx_fig1b). The latter is calculated from a Floquet simulation of either a self-squeezing of the LINC (with an added Kerr to induce $|0\rangle\leftrightarrow|2\rangle$ oscillations), or a full beamsplitting process (see appendix B). It is also simple to derive the relative sensitivity of this process to flux noise, for eq. 6 in the main text:
@@ Appendix B Driven operations with the LINC
To numerically demonstrate such a parametric beamsplitting process, we consider the specific frequency stack of $\omega_{a}=2\pi\times 4.9$GHz, $\omega_{b}=2\pi\times 6.0$GHz, and $\omega_{c}=2\pi\times 6.5$GHz, with Rabi coupling strengths of $g_{ac}=2\pi\times 120$MHz and $g_{bc}=2\pi\times 50$MHz. We choose the LINC circuit parameters as $E_{C}=2\pi\times 100$MHz, $E_{J}=2\pi\times 15.84$GHz, and $E_{L}=2\pi\times 52.8$GHz. With M LINC loops in an array, the system Hamiltonian is given by
@@
Appendix C Evaluating driven performance
We now focus on quantifying the driven coupler’s performance, in the presence of environmental noise and the resonator modes, Alice and Bob. Specifically, we show that the LINC’s is resilient to drive-induced parasitic processes and dressed decoherence, which should lead to an improvement in parametric process fidelity. As a benchmark for comparison, we compare it to a SNAIL at its Kerr-free flux point.
We model the driven SNAIL Hamiltonian in the displaced frame as
@@ Appendix C Evaluating driven performance
where $M$ is the number of SNAIL loops in an array, $N$ is the number of Josephson junctions in the array of each SNAIL loop, and $\theta_{\text{eq,i}}$ is the frustration phase across the i-th branch of the SNAIL loop, as a result of the dc flux penetrating the loop. For consistency, we denoted the phase displacement of the SNAIL as $\phi_{AC}$, even though it has a different physical origin than flux modulation in the LINC circuit. We carefully choose the SNAIL parameters as $E_{C}=2\pi\times 100$ MHz, $E_{J}=2\pi\times 276$ GHz, and $\alpha=0.193$. Under these parameters, the SNAIL circuit yields the same frequency of $\omega_{c}=2\pi\times 6.5$GHz as the LINC, as well as the same $g_{3wm}$, at its Kerr-free point of $\Phi_{\text{ext}}=0.442\Phi_{0}$.
@@
To characterize their performance as microwave beamsplitters, we first examine the rate and resonance condition of the Alice-Bob beamsplitter interaction activated by driven the LINC and SNAIL respectively, as a function of drive amplitude. We employ time-dependent perturbation theory for both couplers in the Floquet mode basis [26] to calculate the beamsplitter rate as a function of coupler state, $|m\rangle$. This amounts to finding the coefficient of the effective Hamiltonian,
$H_{\text{eff}}=g_{\text{BS,m}}\hat{a}^{\dagger}\hat{b}|m\rangle\langle m|+\text{h.c.}$ (33)
@@
When the coupler is in its ground state ($m=0$), this calculation provides the beamsplitter rate reported in the main text. When the coupler is excited to higher levels, we observe state-dependent shifts of the beamsplitter rate and resonance condition (LABEL:appx_fig3a,b). This dispersion makes it clear that if the coupler leaves its ground state and enters a mixed state due to the parametric drive, it can dephase the beamsplitting process. As a result, the drive-induced excitation of the coupler state can lead to process infidelity, with magnitude set by the amount of state-dependent dispersion in $g_{\text{BS}}$ and $\Delta_{\text{BS}}$.
Figure 6: Driven coupler impurity a, Dispersion in beamsplitting strength (left) and resonance condition (right) as a function of coupler (undriven) state $|0\rangle,\dots,|3\rangle$. The LINC (grey) has less dispersion than the SNAIL (yellow), implying that for similar excited state populations, it will have higher process fidelity. b, Steady-state impurity of the driven LINC and the (charge) driven SNAIL, in the presence of decay and flux-noise dephasing, as a function of drive frequency at fixed amplitude ($\phi_{\text{AC}}=0.2\pi$). The simulation also includes additional coupled modes Alice (4.9 GHz) and Bob (6.0 GHz). The LINC remains significantly more pure at all drive frequencies. Dominant processes are explicitly labeled. c, Driven coupler impurity in the presence of two drive tones, each with amplitude $\phi_{AC}=0.1\pi$, and coupled Alice and Bob modes, displaying the effects of parasitic intermodulation products. The two drive frequencies are scaled simultaneously maintaining a fixed ratio of 1:2, to ease Floquet-Markov simulations.
fig]appx_fig3
To study such drive-induced excitation in the driven LINC and SNAIL as well as spurious coupler-resonator interactions, we employ the Floquet-Markov approach. We consider an environment that contains a coupler thermal relaxation channel and a dephasing channel for both circuits. To efficiently capture the effects of additional modes in the system, we model Alice and Bob as two Lorentzians set by their frequencies and linewidths, resulting in a total coupler relaxation spectrum of
$\Gamma_{c}(\omega)$ $=\frac{\omega}{\omega_{q}}(1+n_{B}(\omega,T_{c}))\gamma_{c}$ (34)
$+\sum_{i=A,B}(1+n_{B}(\omega,T_{i}))\frac{g_{i}^{2}}{(\omega-\omega_{i})^{2}+(\frac{\kappa_{i}}{2})^{2}}\kappa_{i}.$
Here, $\gamma_{c}=(26.7\mu\text{s})^{-1}$ is the coupler decay rate, $n_{B}(\omega,T)=1/(e^{\hbar\omega/k_{B}T}-1)$ is the Bose-Einstein occupation factor. We set Alice’s and Bob’s temperature and decay rates to 50mK and $(10\mu s)^{-1}$, higher than the coupler’s, to efficiently capture and visualize the parasitic transitions. We also assume $1/f$ dephasing spectrum for the coupler, $S_{\Phi\Phi}[\omega]=\frac{(1\mu\Phi_{0})^{2}}{\omega}$, with a cut-off frequency of 1 Hz.
Using the Floquet-Markov approach, we examine the impurity of the steady state over different drive frequencies and amplitudes, as a proxy for process infidelity caused by spurious coupler excitation, as well as by other parasitic coupler-resonator interactions. When the two circuits are driven by a single tone (LABEL:appx_fig3b) across the frequency range of 0.5- 2 GHz, both impurities remain small, except for a few peaks that represent coupler-resonator transitions. At higher frequencies above 2 GHz, LINC has substantially lower impurity, benefiting from its parity selection rule that forbids the spurious sub-harmonic excitations that SNAIL encounters at 2.17 GHz and 3.25 GHz. Further, when two drives are simultaneously applied to the circuits (LABEL:appx_fig3c), we observe a significant difference between the LINC and SNAIL, where the former possesses much lower impurities and fewer parasitic transitions than the latter. Finally, when sweeping drive amplitude, such as one might enact in the duration of a pulse sequence, the LINC has negligible drive-induced shifts in its parasitic resonances, as seen in LABEL:fig3c. This is in stark contrast to Kerr-full couplers like a regular transmon, or the DDS.
Overall, LINC’s low driven process dispersion and significantly cleaner, more stable driven spectrum promise an expanded drive space for high-fidelity parametric operations, including when dealing with multiple modes or pumps.
Appendix D Effect of junction and DC flux asymmetries
@@
Figure 7: LINC Kerr in the presence of asymmetries a, Kerr of the LINC mode at $\phi_{\text{DC}}=\pi/2$, computed by direct diagonalization. Black contour lines correspond to $|\alpha_{L}|=250,500,750,\text{and }1000$ KHz respectively. b, Operating point $\phi_{d}$ where the LINC is Kerr-free, as a function of asymmetry. With a minor change in the symmetric DC flux, the Kerr induced by small asymmetry can be nulled. Black contour lines correspond to flux values from $98\%\times\pi/2$ to $102\%\times\pi/2$ in $0.5\%$ increments around the asymmetry-free DC operating point. c, third-order nonlinearity of the LINC dipole mode, $g_{30}$. With finite junction asymmetry, the self-Kerr and cross-Kerr-free points will differ. Contour lines correspond to $|g_{30}|=3,6,9,\text{and }12$ MHz respectively.
@@
$\begin{split}U/E_{L}=&\frac{1}{2}(\hat{\theta}_{c}-2\phi_{\Delta}/3)^{2}\\ +\beta_{\Sigma}\sin(\phi_{AC})&\left[\cos(\phi_{\Delta}/3)\cos(\hat{\theta}_{c})-\sin(\phi_{\Delta}/3)\sin(\hat{\theta}_{c})\right]\\ +\beta_{\Delta}\cos(\phi_{AC})&\left[\cos(\phi_{\Delta}/3)\sin(\hat{\theta}_{c})+\sin(\phi_{\Delta}/3)\cos(\hat{\theta}_{c})\right].\end{split}$ (43)
@@
Figure 8: Driven process strength in the presence of asymmetries a, Fractional strength of the desired $g_{21}$ process in the presence of asymmetries. We see that the strength is primarily dependent upon the asymmetric flux $\phi_{\Delta}$, and is second order with respect to both asymmetries. b, strength of $g_{11}$, the linear coupling of the LINC mode to a symmetric flux drive, as compared to the LINC’s linear coupling to a charge drive $g_{11}^{\text{charge}}\approx E_{L}\theta_{zpf}$. c, Parity protection within the 3rd order nonlinearity. Strength of $g_{12}\phi_{\text{AC}}^{2}$, that would permit subharmonic-driving-like processes, as compared to the desired $g_{21}\phi_{\text{AC}}$, is shown as a function of asymmetry for the drive value $|\phi_{\text{AC}}|=0.2\pi$. The above simulations are performed by numerically computing appropriate derivatives of the potential and $\theta_{zpf}$ as detailed in eqs. 44, 45 and 46. The behavior of these driven process strengths with respect to asymmetry has been separately verified by floquet and/or time-domain simulations.
@@
[…]
@@
Appendix E Effects of parasitic loop inductance
We now analyze the effect that a loop parasitic inductance has on the static nonlinearity of the LINC potential energy. In particular, we assume each Josephson junction in the outer SQUID loop has a stray series inductance ($L_{\text{loop}}$ in LABEL:appx_fig4a), and apply nonlinear current conservation methods developed in [69]. We give intuition for and setup the equations for the analysis below, and then evaluate the equations numerically for actual comparison. Note that because the loop inductance always shares current with an outer junction, it is difficult to directly give analytic solutions that are non-parametric. The case of the series inductance in the coupler arms ($L_{\text{P}}$ in LABEL:appx_fig4a) does not face this issue, and we treat it fully analytically in the next section.